In 4 ABC , AB = AC . Bis (3 ; 6) and C is (7 ; 12). G is the centroid of 4 ABC . G moves on a line x- y = 1. Find coordinates of G

 


 




In



 


4ABC, AB = AC. Bis  (3; 6) and C is (7; 12). G

is the

centroid of



 


4ABC. G moves on a line x-y

= 1. Find coordinates

of



 


G



Grade:10

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear ajinkya

G lies on x-y=1

let co ordinate of G is (h,h-1)

since AB =Ac

co distance og G from B and C will be same

BG = CG

(h-3)2 +(h-1-6)2 = (h-7)2 + (h-1-12)2

h=8

so G (8,7)

 Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin